OpenStudy (anonymous):

whats x ? 15=5^(x+3) -7

OpenStudy (anonymous):

please helppp n_n urgent

OpenStudy (anonymous):

how do i start

OpenStudy (anonymous):

i added 7 to both sides already

OpenStudy (anonymous):

then i am stuck

OpenStudy (vishweshshrimali5):

\[\large{15 = 5^{x+3} - 7}\] \[\large{15+7 = 5^{x+3}}\] \[\large{22 = 5^{x+3}}\] Now, try taking log on both sides.

OpenStudy (anonymous):

log22=log5^x+3 ln22=x+3ln5 x=ln22/ln5 ???? @vishweshshrimali5

OpenStudy (vishweshshrimali5):

No

OpenStudy (vishweshshrimali5):

You were right till this step: \[\large{\ln22 = (x+3)\ln{5}}\] \[\large{\cfrac{\ln22}{\ln{5}} = x+3}\]

OpenStudy (anonymous):

ohhhh subtract three both sides ??

OpenStudy (vishweshshrimali5):

Yeah

OpenStudy (anonymous):

sooo x=ln22/ln5 -3

OpenStudy (vishweshshrimali5):

good

OpenStudy (vishweshshrimali5):

yes

OpenStudy (anonymous):

thankkkk you @vishweshshrimali5

OpenStudy (vishweshshrimali5):

No problem.